How many electrons per second pass a point in a circuit when a current of 1.60A is established in the circuit? (Charge on an electron = 1.6x 10^-19 C)



1
Expert's answer
2021-05-06T16:50:24-0400

The information given is

The charge of the electron is "|e^{-}|= 1.6x 10^{-19}\;C"

The electric current is "I=1.60\;\text{A}"


The electric current is equal to the number of electric charges that pass through a wire in a certain time, that is to say.


"I=\dfrac{Q}{t}"


Where.

"Q" is the electric charge.

"t" is the time.


Obtaining the expression for the electric charge.


"I=\dfrac{Q}{t}\\
\dfrac{Q}{t}=I\\
Q=I\;t"


Evaluating numerically for t = 1 s


"Q=I\;t\\
Q=1.60\;\text{A}\times 1\;\text{s}\\
Q=1.60\;\text{C}"


The total electrical charge that passes is "Q=1.60\;\text{C}"


Now, the number of electrons is given by

"N=\dfrac{Q}{|e^{-}|}"

Where.

"Q" is the electric charge.

"|e^{-}|" is the magnitude of the electric charge of an electron.


Evaluating numerically.


"N=\dfrac{Q}{|e^{-}|}\\
N=\dfrac{1.60\;\text{C}}{1.6x 10^{-19}\;C}\\
\\
N=1\times 10^{19}\;\text{Electrons }"


Finally.

The number of electrons is "\displaystyle \color{red}{\boxed{1\times 10^{19}\;\text{Electrons }}}"